# Fairly Accurate: Fairness-aware Multi-group Target Detection in Online Discussion

### v2

@@ Abstract
Target-group detection is the task of detecting which group(s) a social media post is “directed at or about”, with various applications, such as targeted-marketing. In this work, we focus on the fairness implications of target-group detection in the context of toxicity detection, where the perceived harm of a post often depends on which group(s) it targets. Because toxicity is highly contextual, language that appears benign in general may be harmful when targeting specific demographic groups. It is thus important to first detect which group(s) are being targeted by a post as a precursor to the subsequent task of determining whether the post is toxic given the group(s). Target-group detection is also challenging: a single post may simultaneously target one to many groups, and we must detect groups fairly in order to promote equitable treatment. We show that our proposed approach to fairness-aware multi target-group detection not only reduces bias across groups, but also achieves competitive predictive performance, outperforming existing fairness-aware baselines. To spur future research on fairness-aware target-group detection and support competitive benchmarking, we also share our code.
@@ 1 Introduction
In online discussions, it is important to understand which groups(s) are being referred to in order to properly interpret the discourse. Target-group detection is the task of detecting which group(s) a social media post is “directed at or about” [38], with downstream applications such as targeted-marketing and toxicity detection. In the latter case of toxicity detection, one asks whether a post is hateful or discriminatory toward specific demographic groups (vs. a task such as offensive language detection [12, 49], where offensiveness does not involve specific demographics). Toxicity is also often highly contextual, with language that might seem benign taking on harmful undertones when it targets specific groups [11, 40, 29, 24, 6]. Target-group detection thus represents an important, distinct, and separable task from any downstream tasks based on detected targets.
A key aspect of target-group detection is that a single post may simultaneously target one to many groups. However, prior work has predominantly framed target-group detection as a single-label task, assuming each post targets at most one group [5, 40]. Moreover, most toxicity detection datasets [47, 12] largely lack granular demographic annotations or focus narrowly on binary classifications, failing to account for the complex interplay of cultural, linguistic, and identity-based differences in what constitutes harmful content. Simplifying the task in this way represents a natural starting point for research but misses the multi-label nature of real-world discourse, where harmful language can implicate multiple groups simultaneously.
Another important consideration is that it is not enough to detect groups accurately; we must also detect groups fairly in order to promote equitable treatment across diverse user demographics. For example, in targeted-marketing of appropriate pharmaceutical drugs to patient populations with different diseases, both groups are equally important to detect because each deserves the same opportunity to benefit from the appropriate drug. Detection models trained without fairness constraints may underperform on marginalized groups, reinforcing existing societal biases or harms and leading to disproportionate impact in downstream applications.
@@
Table 1: Sample comments from the MHS Corpus [38], showing the multi-label format for groups. Note that groups can be targeted by a post regardless of the post’s toxicity. To minimize harm and provide a content warning to readers, we have partially redacted toxic language in this table (e.g., profanity masked with asterisks or ellipses).
To date, most fairness literature has focused on decision-making scenarios for allocating goods or burdens, such as college admissions or bail decisions, where false positives and false negatives have asymmetric costs [26]. Target-group detection is notably different. With the pharmaceutical drug example above, misidentifying either group would be equally harmful since either would miss the opportunity to benefit from the appropriate drug for their disease. This motivates symmetric error weighting, promoting equitable treatment of groups and mitigating biases that could amplify disparate impact.
@@
When considering target-group detection for toxicity detection, symmetric error weighing is also desirable. Toxicity detection itself can be viewed as a task of allocating goods and burdens: from the perspective of targets, content moderation that protects the target from toxic content is a good, and therefore gaps in false negatives in toxicity detection may constitute an issue of distributive justice [35, 20]. In contrast, target-group detection is a multi-label task wherein labels are demographic groups themselves (e.g., Latinx, Black), and both error types, misidentifying one group for the other or vice-versa, are equally harmful. For instance, mistaking a post targeting Latinx as one targeting Black is as undesirable as the reverse.
In this work on fairness-aware target-group detection, we address the two key aspects above: that a single post may target one to many groups (e.g., see Table 1); and that detection must be conducted fairly to promote equitable treatment of groups. We use the recent, large-scale Measuring Hate Speech (MHS) Corpus [38] that labels granular information about the type of content in a post (e.g., inciting violence, humiliating, etc.), as well as identifying all groups targeted by each post. Crucially, note that targets are identified for both toxic and non-toxic posts; a social media post can be “directed at or about” [38] one or more target group(s) without being toxic. This dataset lets us to move beyond simplistic definitions of targeted groups by encompassing diverse demographic dimensions: race, gender, religion, and other axes of identity.
@@
In regard to the symmetric cost of group-level errors in target-group detection, Accuracy Parity (AP) [54] provides a natural fairness criterion in seeking balanced predictive accuracy across groups to mitigate bias. To optimize for AP during model training, we adopt and extend the recently proposed Group Accuracy Parity (GAP) loss [19], which penalizes accuracy deviations across groups. Specifically, because the existing GAP formulation is limited to binary settings and does not scale to multi-label scenarios, we introduce $GAP_{multi}$, a scalable loss function that extends GAP to enable fair detection when posts target multiple groups. Our formulation enables fair optimization in realistic multi-group settings while preserving the convergence behavior of the original GAP formulation, thereby ensuring group fairness and global utility.
@@
While Equalized Odds [21] could be considered as an alternative to AP to independently enforce equality in both error types, EO does not actually treat both types of errors equally, and its use in target detection setting could lead to unintended consequences. Through a formal theoretical analysis, we present an impossibility result: under realistic scenarios (e.g., unequal group base rates), one cannot simultaneously satisfy both AP and EO. Moreover, we show that enforcing EO can actually degrade performance for some groups. These findings reinforce AP as a more appropriate fairness objective for our task.
@@
We evaluate multi-group $GAP_{multi}$ loss against a set of baseline loss functions, including traditional fairness-aware losses and standard optimization strategies. We assess two key objectives: balanced performance across demographic groups (group-fairness) and overall predictive accuracy (utility). $GAP_{multi}$ not only achieves better group-fairness (by minimizing performance disparities impact across groups), but also maintains competitive overall accuracy. Our work thus effectively addresses the unique challenges of multi target-group detection, consistently reducing bias across groups while maintaining similar overall utility.
@@
To summarize, we propose a fairness-aware target-group detection framework that optimizes for both utility and group fairness in a multi-label setting. Jointly optimizing for group fairness and utility supports making predictions that are both fair and accurate, addressing challenges posed by imbalanced datasets while capturing group-specific biases. Our work bridges theoretical fairness objectives with real-world applicability, offering a robust and practical solution for group-identification tasks. To spur future research on fairness-aware target-group detection and support competitive benchmarking, we also share our code [^1] URL: https://github.com/smjtgupta/GAP.
@@ 2 Related Work
Fairness in machine learning has been extensively studied across many critical applications such as risk assessment [21], college admissions [26], loan approval [9], and recidivism [2]. These domains typically involve tasks where asymmetric errors, i.e., allocating goods or burdens, are central to the decision-making process. For instance, in loan approvals, false negatives (denying credit to a qualified individual) may carry a greater societal cost than false positives (approving a loan for someone who defaults). Similarly, in recidivism prediction, false positives (predicting someone will reoffend when they would not) disproportionately impacts marginalized groups, making error prioritization critical.
@@
For this reason, fairness measures such as Equalized Odds [21] and Demographic Parity [53] are popular. Due to their focus on error rate alignment across groups, these measures are particularly suited for tasks where the interpretation of error rates is tied to the societal implications of unequal access to resources or unjust burdens. In contrast, we investigate multi target-group detection, which presents a fairness challenge where error types are symmetric: misidentifying a targeted group (false negative) and incorrectly assigning a group (false positive) are equally undesirable. This symmetry is relatively rare in traditional fairness applications but critical here, motivating the use of Accuracy Parity (AP) [54] as a natural and principled measure. While AP has been studied conceptually [3, 10, 32], relatively little work has explored its role in multi-group detection settings.
Target-group detection in Toxicity modeling.
@@
Identifying which demographic group(s) are targeted by toxic language has become a crucial component in modern toxicity detection pipelines. Early work by Blodgett et al. [5] framed target-group detection as a single-label classification task across four groups, selecting the most likely group during inference. Similarly, many datasets [12, 16] treat group identification as binary or mutually exclusive, assuming each post targets only one group. However, real-world toxic content often implicates multiple communities simultaneously. While recent studies discuss the need the considering group-targeted toxicity [45] or seek to model it [28, 18], to the best of our knowledge, all have assumed oracle detection of target groups based on annotated corpora.
Fairness-aware loss functions.
@@
A central challenge in fairness-aware machine learning is aligning the training objective (i.e., the loss function) with the fairness criterion used for evaluation. When this alignment is weak, as in the case of many surrogate loss functions [50, 51, 7], models may experience metric divergence, especially under sensitive hyperparameter regimes [33, 31, 52, 44]. These surrogate losses typically optimize for proxies (e.g., cross-entropy with reweighting) rather than directly minimizing fairness violations. In contrast, strictly-mapped loss functions [19, 42] are constructed to directly correspond to fairness measures. These formulations offer a smoother and more reliable optimization path — as the loss decreases, the corresponding fairness metric improves predictably.
Datasets for Target-Group Detection.
The development of fairness-aware target-group detection models has been hindered by the lack of datasets that provide fine-grained, multi-label annotations of demographic groups targeted in social media posts. Early toxicity datasets [47, 12] label content with binary hate labels and limited demographic group references. Subsequent corpora, including those by Blodgett et al. [5], Founta et al. [17] similarly fail to represent the complex overlap of targeted identities present in real-world discourse. Moreover, few datasets enable robust fairness evaluation, as they lack annotator consensus indicators, group-level prevalence statistics, or non-toxic examples that still target identity groups. To address these gaps, we adopt the MHS Corpus [38], a large-scale dataset that explicitly annotates several demographic groups and captures both toxic and non-toxic instances of group references. The dataset supports multi-label annotations per post making it uniquely suitable.
3 Multi Target-group Detection
@@
Group-specific toxicity often manifests through distinct language patterns: coded language may be used against certain ethnicities and innocuous terms can become racialized slurs in certain contexts. Certain forms of toxicity such as gender-based hostility often also involve stereotype-driven insults or harassment. Such content, though superficially benign, may acquire harmful undertones when directed at particular communities. Without explicitly detecting the intended target group(s), toxicity detection models may risk generalizing prediction, leading to misclassifications and disparate impact for the statistical minority groups [40, 11].
Use of annotator-provided target labels enables training of group-conditioned toxicity detection models [18, 15, 20], capturing nuanced differences between shared and group-specific toxicity patterns. However, deploying such systems in practice requires solving the real-world task of target group detection in the wild. Unlike controlled annotation settings, real-world content typically lacks annotated labels identifying target groups; user reports may provide this information, but proactive toxicity detection prompts the need for an accurate and fair target-detection module.
3.1 Problem Statement
To formally define the fairness-aware multi target-group detection task, we assume a dataset $X\in\mathbb{R}^{N\times F}$, where each of the $N$ samples represent a post, and $F$ denotes the dimensionality of extracted features. Let $G$ be the total number of demographic groups. Let $Y\in\mathbb{R}^{N\times G}$ denote the label space, where $y=\{0,1\}^{G}$ represents the set of $G$ binary labels, indicating whether each group $g\in G$ is targeted ($1$) or not (00) by the post $x$. This leads to a multi-label classification problem setting [30, 23], since a post $x$ can target multiple groups simultaneously. The objective is to learn a prediction function set $f:X\rightarrow\{0,1\}^{G}$ that maps each post $x$ to its corresponding multi-group label $y$. For each group $g$, we define the group-specific and overall model accuracy as $Acc_{g}$ and $Acc_{overall}$ respectively.
$Acc_{g}$ $=\frac{\textrm{Number of correct predictions for group }g}{\textrm{Total number of predictions for group }g}$ (1)
$Acc_{overall}$ $=\frac{\sum_{g=1}^{G}\textrm{Number of correct predictions for group }g}{\textrm{Total number of predictions over all groups }G}$ (2)
@@ 3.1 Problem Statement
From an overall utility perspective, we aim to maximize overall accuracy across all groups $G$. However, to ensure fairness, we also seek to minimize disparities in predictive accuracy across groups. We define the inter-group disparity as the deviation of each group’s accuracy from the mean overall accuracy. Thus, we have the following set of optimization problems in Eq. 3 and Eq. 4:
[…]
@@
To jointly optimize for both utility and fairness, we adopt a Single Objective Optimization (SOO) goal, combining these objectives, Eq. 3 as primary and Eq. 4 as secondary, with a regularization term $\lambda\in(0,\infty)$ in Eq. 5. This formulation explicitly balances model utility with group-level fairness, penalizing uneven performance across groups.
[…]
@@ 4 From Measure to Optimizer: AP & GAP
Accuracy Parity (AP)
To address the trade-off between accuracy vs. group fairness, AP measures the degree to which predictive accuracy varies across demographic groups. It is suitable for tasks with symmetric error costs, such as target-group detection. AP is well-studied under various names: Accuracy Equity [13], Equality of Accuracy [22], Accuracy Parity [54], Equal Accuracy [32], Overall Accuracy Equality [3] and Accuracy Difference [10].
To date, AP has predominantly been discussed as an evaluation measure, while less effort has been devoted to developing suitable loss functions. This may stem in part from the fact that a bulk of algorithmic fairness literature has been devoted to the use of ML in the allocation of goods and burdens, where this is typically not deemed a desirable fairness measure since it does not capture the different costs of different types of errors. For multi target-group detection, AP is particularly well-suited because it aligns with the symmetric error costs inherent in the task. In this setting (as defined in Eq. 4), errors from all groups are treated equally, irrespective of group membership. For instance, misclassifying a post targeting one group as targeting another group is equally undesirable, regardless of the directionality of the error. A particular attribute of the target-group detection task also merits note: the group and the label are the same. AP thus measures the extent to which all groups are treated equitably.
Group Accuracy Parity (GAP)
To directly optimize for Accuracy Parity (AP), we adopt Group Accuracy Parity (GAP) loss [19], the first differentiable loss function for AP that we are aware of. Unlike surrogate losses that loosely correlate with fairness metrics, GAP strives to minimize the loss directly to improve AP during training. The GAP formulation shown in Eq. 6 combines the overall error (OE) with a regularization term measuring the disparities in cross entropy (CE) error across the binary groups ($g=0,1$). It allows different weighted variants of accuracy w.r.t. chosen entropy, depending on the evaluation need.
@@
The regularization term evaluates to zero when the errors across the two groups are same, i.e., perfect AP score of 0. The squared two-norm formulation supports a smooth loss surface that results in better optimization trajectory. While formulated for binary group settings, this structure forms the foundation for our multi-label extension, as discussed next.

### v3

# Fairness-Aware Multi-Group Target Detection in Online Discussion

@@ Abstract
Target-group detection is the task of detecting which group(s) a piece of content is “directed at or about”. Applications include targeted marketing, content recommendation, and group-specific content assessment. Key challenges include: 1) that a single post may target multiple groups; and 2) ensuring consistent detection accuracy across groups for fairness. In this work, we investigate fairness implications of target-group detection in the context of toxicity detection, where the perceived harm of a social media post often depends on which group(s) it targets. Because toxicity is highly contextual, language that appears benign in general can be harmful when targeting specific demographic groups. We show our fairness-aware multi-group target detection approach both reduces bias across groups and shows strong predictive performance, surpassing existing fairness-aware baselines. To enable reproducibility and spur future work, we share our code online [^1] URL to be shared upon paper acceptance..
@@ 1 Introduction
What is target-group detection? In online discussions, properly interpreting the discourse and managing content requires knowing which groups(s) are being referred to in discussion. Target-group detection is the task of detecting which group(s) a piece of content is “directed at or about” [35]. Applications include targeted marketing, content recommendation, and group-specific content assessment. In the first two cases, the group that may deem certain content relevant is the target. For content assessment, the group that a post speaks to or about is the target.
Why does it matter? In fact-checking, identifying a post’s target is key to assessing the post’s checkworthiness (i.e., appropriately prioritizing it for fact-checking) since “potential harm to specific groups" can be one of the criteria used for prioritization [33, 26]. In toxicity detection, whether a post is hateful or discriminatory is often highly contextual, with language that might seem benign taking on harmful undertones if the target is a specific group [10, 37, 28, 22, 5]. Target-group detection is thus important, distinct, and separable from any downstream tasks based on detected targets, such as assessing if a post is toxic [19] depending on the target(s).
What are the challenges? First, a single post may target multiple groups. If we consider toxicity detection, prior work has largely framed target-group detection as a single-label task, assuming each post targets at most one group [4, 37]. Most toxicity datasets [40, 11] also lack granular demographic annotations or focus narrowly on binary classifications. This fails to account for the complex interplay of cultural, linguistic, and identity-based differences in what constitutes harmful content. It also misses the multi-label nature of real-world discourse, where language can implicate multiple groups simultaneously, as illustrated below in Table 1.
Second, it is not enough to detect groups accurately; we must also detect groups fairly for equitable treatment across diverse demographics. For example, in personalized content recommendation, such as suggesting movies or music to diverse user groups, accuracy disparities in target group detection could result in one or more groups failing to receive content that is relevant to them. In toxicity detection, disparities in target group detection can have downstream detrimental impacts on toxicity detection models that assess toxicity conditional on the target group [19, 14].
Table 1: Sample comments from the MHS Corpus [35], showing the multi-label format for groups. Note that groups can be targeted by a post regardless of the post’s toxicity. We have partially redacted toxic language in examples shown.
@@
The value of symmetric error weighting. To date, most fairness literature has focused on scenarios for allocating goods or burdens (e.g., college admissions or bail decisions) wherein false positives and negatives have asymmetric costs [24]. Target-group detection is different. Consider content recommendation, where preferences (e.g., in movies or music) vary across groups. If our goal is to promote fairness across all groups, there is no universal notion of “positive" and “negative" class because content relevant to one group may be irrelevant to another, and all groups should be equally likely to encounter relevant content. Symmetric error weighting across groups treats groups equitably.
@@
The same principle holds in toxicity detection, which can also be viewed as a task of allocating goods and burdens. From the perspective of targets, content moderation that protects the target from toxic content is a good, and so gaps in false negatives in toxicity detection may constitute an issue of distributive justice [33, 19]. In contrast, the upstream task of target-group detection is a multi-label task: labels are demographic groups themselves (e.g., Latinx, Black), and errors misidentifying one group for the other are equally undesirable. For instance, mistaking a post targeting Latinx as one targeting Black is equally undesirable as the reverse.
@@
How can we optimize for symmetric errors across groups? The Accuracy Parity (AP) measure [45] provides a natural fairness criterion for the task of multi-target group detection. To optimize for AP during model training, we extend the Group Accuracy Parity (GAP) loss function [18], which penalizes accuracy deviations across groups. Specifically, because GAP is limited to binary settings, we introduce $GAP_{multi}$, a scalable loss that extends GAP to enable fair detection when posts target multiple groups. Our formulation enables fair optimization in realistic multi-group settings while preserving the convergence behavior of the orginal GAP loss.
@@
Why not use Equalized Odds (EO)? While EO might be considered instead of AP to independently enforce equality in both error types, we show that EO [20] does not actually treat both types of errors equally, and as a consequence, gives priority to some groups over others. Through a formal theoretical analysis, we present an impossibility result: under realistic scenarios (e.g., unequal group base rates), one cannot simultaneously satisfy both AP and EO. Moreover, we empirically show that enforcing EO can actually degrade performance for some groups. Consequently, its use in target detection setting could lead to unintended consequences.
Evaluation context: toxicity detection. Having motivated fairness-aware multi-group target detection across several domains, we now focus our empirical study in the context of toxicity detection. Note that we do not evaluate toxicity detection itself in this study. We only evaluate the preceding task of multi-group target detection. As discussed earlier (paragraph 2), identifying the target of a post is a necessary precursor to assessing if a post is toxic or not.
We use the large-scale Measuring Hate Speech (MHS) Corpus [35] and HateXplain [29] datasets for our study because these: 1) include and annotate posts targeting multiple groups; 2) annotate targets for both toxic and non-toxic posts (a post can be “directed at or about” (Ibid.) one or more target group(s) without being toxic); and 3) include demographic group annotations: race, gender, religion, among others. Evaluation on these two datasets, which include data from multiple platforms (Twitter, Reddit, YouTube, Gab), validates generalization across different discourse styles and moderation environments, testing robustness beyond platform-specific artifacts.
@@
We compare our proposed multi-group $GAP_{multi}$ loss to several baseline loss functions, including traditional fairness-aware losses and standard optimization strategies. We assess two key objectives: group-fairness (balanced performance across demographic groups) and utility (overall predictive accuracy). $GAP_{multi}$ not only achieves better group-fairness (by minimizing performance disparities impact across groups), but also maintains competitive overall accuracy. Our work thus effectively addresses the unique challenges of multi-group target detection, consistently reducing bias across groups while maintaining similar utility.
@@
Summary of contributions: We motivate and propose a fairness-aware multi-group target detection framework that jointly optimizes overall utility and group fairness in a multi-label setting. The proposed $GAP_{multi}$ loss enables parallel GPU computations of group-pairwise errors, allowing a constant factor scaling to an arbitrary number of groups. This supports scalable optimization needed for deployment in dynamic online platforms. Empirical validation across datasets demonstrate the robustness of our loss function, bridging theoretical fairness objectives with real-world applicability. This provides a robust and practical solution for group-identification tasks by making predictions both fairly and accurately. To spur future research and support benchmarking, we also share our code online (see abstract).
@@ 2 Related Work
Fairness in Machine Learning has been studied across critical applications such as risk assessment [20], college admissions [24], loan approval [8], and recidivism [1]. These domains typically involve tasks where asymmetric errors, i.e., allocating goods or burdens, are central to the decision-making process. For instance, in loan approvals, false negatives (denying credit to a qualified individual) may carry a greater societal cost than false positives (approving a loan for someone who defaults). Similarly, in recidivism prediction, false positives (predicting someone will reoffend when they would not) disproportionately impacts marginalized groups, making error prioritization critical. For this reason, fairness measures such as Equalized Odds [20] and Demographic Parity [44] are often employed. Given their focus on error rate alignment across groups, these measures are well-suited for tasks where the interpretation of error rates is tied to the societal implications of unequal access to resources or unjust burdens.
@@
In contrast, multi-group target detection offers a different fairness challenge wherein error types are symmetric: misidentifying a targeted group (false negative) and incorrectly assigning a group (false positive) are equally undesirable. Such symmetry is relatively rare in classic fairness applications but critical here, motivating our focus on Accuracy Parity (AP) [45] as a natural and principled measure. While AP has been studied conceptually [2, 9, 31], relatively little work has explored it for multi-group detection.
@@
In toxicity modeling, identifying which demographic group(s) are targeted by toxic language has become a crucial component in modern toxicity detection pipelines. Early work by Blodgett et al. [4] framed target-group detection as a single-label classification task across four groups, selecting the most likely group during inference. Similarly, many datasets [11, 15] treat group identification as binary or mutually exclusive, assuming each post targets only one group. However, real-world toxic content often implicates multiple communities simultaneously. While recent studies discuss the need the considering group-targeted toxicity [39] or seek to model it [25, 17], to the best of our knowledge, all have assumed oracle detection of target groups based on annotated corpora.
@@
A central challenge in fairness-aware machine learning is aligning the training objective (i.e., the loss function) with the fairness criterion used for evaluation. When this alignment is weak, as in the case of many surrogate loss functions [42, 43, 6], models may experience metric divergence, especially under sensitive hyperparameter regimes [32]. These surrogate losses typically optimize for proxies (e.g., cross-entropy with reweighting) rather than directly minimizing fairness violations. In contrast, strictly-mapped loss functions [18, 38] are constructed to directly correspond to fairness measures. These formulations offer a smoother and more reliable optimization path — as the loss decreases, the corresponding fairness metric improves predictably.
@@
Group-specific toxicity often manifests through distinct language patterns: coded language targeting certain ethnicities or innocuous terms as racialized slurs in context. Gender-based hostility often involves stereotype-driven insults, where superficially benign content, may acquire harmful undertones when directed at particular groups. Without detecting the intended target groups, toxicity detection models risk generalization, causing misclassifications and disparate impact for minority groups [37, 10]. Target labels from annotators enable group-conditioned toxicity detection models [17, 14, 19], that capture nuances between shared and group-specific toxicity patterns. However, real-world deployment requires detecting target groups in-the-wild, as content lacks pre-annotated labels. While user reports may provide this data, proactive detection demands an accurate and fair target-detection module in place.
3 Multi-Group Target Detection
Problem Statement. Given a collection of $N$ posts, we represent each post as a feature vector in $\mathbb{R}^{F}$ using a pre-trained text encoder, yielding a dataset $X\in\mathbb{R}^{N\times F}$. For a set of demographic groups $G$ (e.g., race, religion, gender), let the label space be $Y\in\mathbb{R}^{N\times|G|}$. Each post $x_{i}$ has a corresponding multi-hot label vector $y_{i}=\{0,1\}^{|G|}$, where $y_{i}[g]=1$ indicates that group $g\in G$ is targeted ($1$) or not (0) by post $x_{i}$. This yields a multi-label classification problem [30, 21] where a single post $x_{i}$ can target multiple groups. We seek to learn a prediction function $f:X\rightarrow\{0,1\}^{|G|}$ mapping each post $x_{i}$ to its multi-group associated label $y_{i}$.
@@ 3 Multi-Group Target Detection
We report group-specific error $err_{g}$ for each group $g$, and overall (macro-average) error over all groups, $\overline{err}_{overall}$. From a utility perspective, we aim to minimize this $\overline{err}_{overall}$. However, to ensure fairness, we also seek to minimize disparities in predictive performance across groups. We define the inter-group disparity ($dev_{overall}$) as the summation of deviation of each group’s error from the mean overall error for predicted labels $\hat{Y}$. Thus, we have the following set of optimization problems in Eq. 1 and Eq. 2:
$\underset{\hat{Y}}{\min}\,\,\overline{err}_{overall}=\frac{\sum_{g=1}^{|G|}err_{g}}{|G|}$ (1)
$\underset{\hat{Y}}{\min}\,\,dev_{overall}=\sum_{g=1}^{|G|}|err_{g}-\overline{err}_{overall}|$ (2)
@@
To jointly optimize for both utility and group fairness, we adopt a Single Objective Optimization (SOO) goal, combining these objectives, Eq. 1 as primary and Eq. 2 as secondary, with a regularization term $\lambda\in(0,\infty)$ in Eq. 3. This formulation explicitly balances model utility with group-level fairness, penalizing uneven performance across groups.
$\text{{SOO goal:}}\qquad\underset{\hat{Y}}{\min}\,\,(\overline{err}_{overall}+\lambda\cdot dev_{overall})$ (3)
@@ 4 From Measure to Optimizer: AP & GAP
Accuracy Parity (AP). To balance accuracy vs. group fairness, AP measures the degree to which predictive accuracy varies across demographic groups. As discussed in Section 1, AP is well-aligned with multi-group target detection because it aligns with the symmetric error costs inherent in the task, treating errors from all groups equally, irrespective of group membership (Eq. 2). For instance, misclassifying a post targeting one group as targeting another group is equally undesirable, regardless of the directionality of the error. A particular attribute of the target-group detection task also merits note: the group and the label are the same.
Group Accuracy Parity (GAP). To directly optimize for Accuracy Parity (AP), we adopt Group Accuracy Parity (GAP) loss [18], the only differentiable loss function for AP that we are aware of. Unlike surrogate losses that loosely correlate with fairness metrics, GAP strives to minimize the loss directly to improve AP during training. The GAP formulation shown in Eq. 4 combines the overall error (OE) with a regularization term measuring the disparities in cross entropy (CE) error across binary groups ($g=0,1$). It allows different weighted variants of accuracy w.r.t. chosen entropy, depending on the evaluation need.
@@
The regularization term evaluates to zero when errors across the two groups are same, i.e., perfect AP score of 0. The squared two-norm formulation supports a smooth loss surface resulting in better optimization trajectory. While formulated for binary group setting, this structure forms the foundation for our multi-label extension, as discussed next.